at a certain time of a day a vertical yardstick casts a 20 inch shadow.how tall in feet is a flafpole that casts a shadow of 15 feet at the same time?
a)20
b)22
c)25
d)27
e)30

Answer by jim_thompson5910(35256)

3/20 = x/(15*12)

3/20 = x/180

3*180 = 20x

540 = 20x

540/20 = x

x = 27

So the answer is 27 ft.